OpenStudy (anonymous):

Y=x^2 +10x+c;(-5,-27)

OpenStudy (anonymous):

\[Y=x^2 +10x+c;(-5,-27) \]

OpenStudy (anonymous):

are you supposed to find c?

OpenStudy (anonymous):

replace x by -5, set the result = -27 and that should work yes?

OpenStudy (anonymous):

\[(-5)^2+10\times (-5)+c=-27\] \[25-50+c=-27\] \[-25+c=-27\] \[c=-27+25\] \[c=-2\]